# Dynamical correlations and nonequilibrium sum rules in photodoped Hubbard ladders

### v2

@@ IV.3 Charge dynamical structure factor
In contrast to the spin sector, the emergence due to pumping of an intraband charge response at low energies $\omega\sim 0$ is unambiguous because of the large Mott gap. This intraband response, due to scattering from pump induced doublons, has a pronounced dispersion and becomes gapless on approaching $\mathbf{k}=(0,0)$. After pumping the intraband charge response is a strong feature, with more spectral weight than the interband charge response in the original equilibrium system. For the $A_{0}=1$ leg pump the ratio of the integrated charge response (including the suppressed static peak at $\mathbf{k}=(0,0)$ and $\omega\sim 0$) after pumping to that in equilibrium is $W_{n}(t=10)/W_{n}(t=-10)\approx 1.2$. For the $A_{0}=1$ rung pump the same ratio gives $\approx 1.1$.
@@
Figure 13: Charge correlations with $\mathbf{k}=(\pi,\pi)$, at $t=10$ after pumping along the legs ()($\parallel$)( ∥ ) or rungs ()perpendicular-to($\perp$)( ⟂ ).
@@
Figure 14: Charge correlations with $\mathbf{k}=(0,\pi)$, at $t=10$ after pumping along the legs ()($\parallel$)( ∥ ) or rungs ()perpendicular-to($\perp$)( ⟂ ).
@@
As for the spin response, the detailed effects of the different pumps are best seen by considering cuts at fixed wave vectors. Figs. 13 and 14 show cuts through the charge response at fixed $\mathbf{k}$. Similar to the spin dynamics, the equilibrium charge response at both $(\pi,\pi)$ and at $(0,\pi)$ appears as a resolution limited peak (except for a small feature close to $\omega=6$ for $k_{y}=\pi$). Pumping suppresses this peak and makes it asymmetric, as well as producing a strong low-energy, intraband response. The energetically similar $A_{0}=0.45$ leg direction, and $A_{0}=1$ rung direction pumps lead to a very similar suppression of the interband response at $\omega\sim 10$. However, there is pump anisotropy visible in the intraband response that develops. Figs. 13 and 14 show that the rung pump produces a weaker response at $\mathbf{k}=(0,\pi)$ than at $\mathbf{k}=(\pi,\pi)$ with an asymmetric energy profile that differs qualitatively between the two wave vectors. In comparison, the $A_{0}=0.45$ leg direction pump produces a very similar post pump response in terms of intensity and shape at both $k_{y}=\pi$ and $k_{y}=0$.
IV.4 Comparison with the sum rule
@@
We finish this section by checking that the integrated spin and charge responses obey Eq. 12. Table 2 displays the results of evaluating $W_{s}$ and $W_{c}$ numerically using two different methods, at three different times: $t=-10$ before the pump, $t=10$ after the pump, and $t=\pi/2$ half way through the pump. The first method involves evaluating Eq. 9 and Eq. 11 using iTEBD data for $\langle S^{2}_{i,l}(t)\rangle$ and $\langle d_{i,l}(t)\rangle$. This method confirms the accuracy of the iTEBD results by satisfying the sum rule $4W_{s}+W_{c}=2G\rho$ to high precision ($<10^{-9}$ missing weight).
@@
The second method we use is to integrate the dynamical response $I(\mathbf{k},\omega,t)$, determined using TEBD for the 32 site ladder, over $\omega$ and $\mathbf{k}$. The results of this method for $W_{s}$ and $W_{c}$ are expected to differ from those for iTEBD due to finite size effects and errors introduced by the single integral approximation, Eq. 6. The latter can also cause departures from the sum rule. Concomitantly, in Table 2 the worst agreement with the sum rule occurs at $t=\pi/2$ where the single integral approximation is least accurate. However, the largest error, seen for the leg direction $A_{0}=0.45$ pump, is still small and $<4\%$.
@@ V Conclusions
We have studied the nonequilibrium dynamics of the half-filled Hubbard ladder when it is pumped by an idealised optical pulse, using MPS techniques. We find that the antiferromagnetic correlations that dominate the equilibrium state are significantly reduced in the post pump state: relative to equilibrium AF correlations are strongly suppressed at nearest and next nearest neighbour separations, before exponential decay sets in with a slightly reduced correlation length. Charge correlations in the initial, equilibrium, state are very weak, but they are strongly enhanced by the pump until they rival the magnetic correlations. We have shown that the anisotropic nature of the ladder means that the energy imparted by the pump is dependent on the pump direction. In general, pumping with a polarisation that alters the hopping parameter along the legs of the ladder leads to greater energy transfer and disruption of spin correlations than pumping polarised in the rung direction. Leg direction pumping also leads to saturation of the transferred energy at lower fluence than for rung direction pumping.
@@
We interpret our results in terms of the hard-core nature of the doublon charge excitations of the system. Photodoping excites electrons from the lower to upper Hubbard band, creating doublons. As they are hard-core, each doublon reduces the available phase space for the creation of new doublons, and this leads to a reduced interband response at energies $\omega\sim U$. For pumping near saturation the interband response is strongly suppressed. The photodoping induced doublons lead to a new low energy response due to intraband transitions. This response is sharply defined and dispersing. In spin systems the corresponding intraband behaviour is known as a Villain mode[40]. A similar intraband charge response has been noted in the equilibrium Hubbard chain at temperatures that are finite, but small compared to the Mott gap [41]. A consequence of the strongly correlated nature of excitations in gapped, hard-core, spin systems is that the thermal broadening of the interband response is strongly asymmetric[38, 15, 16]. We see similar asymmetry in the interband spin and charge responses after pumping.
@@
We have not found a robust value of $\beta^{-1}$ that is shared for both spin and charge structure factors and at different wave vectors (for details see appendix C). We conclude that, like the chain, the post pump state cannot be simply described using a thermal distribution. Indeed Ref. [22] has used a generalised Gibbs ensemble (GGE) approach to characterise the optically pumped Hubbard chain at late times, with the conservation of doublon number leading to an additional generalised temperature. Checking if the post pump Hubbard ladder can be described using a GGE would be an interesting subject for future work.
@@ Acknowledgements.
The authors thank Andreas Weichselbaum and Mark Dean for useful discussions. RMK was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, under Contract No. DE-SC0012704. This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/L010623/1].
Appendix A Effect of bond dimension
@@
Our simulations are carried out using infinite and finite MPS with bond dimensions of up to $\chi=3000$, and we perform time evolution using iTEBD and TEBD with a second order Trotter decomposition and time step, $\Delta t=0.01$. In general we find truncation errors of order $10^{-6}$ at the latest times we report, but because correlations become very short ranged after the pump this does not significantly affect our results (especially once the experimental resolution $\sigma_{X}$ is taken into account). As an example, in Fig. 15 we plot $\langle d_{i,l}\rangle$ for the rung direction pump with $A_{0}=1.0$ using iTEBD with different bond dimensions, $\chi$.
@@
Figure 15: Doublon and holon density as the Hubbard ladder is pumped in the rung direction, i.e. $(\perp,A_{0}=1.0)$, at different bond dimensions $\chi$.
@@
The rung pump case breaks parity symmetry between the legs leading to a less compact structure for the MPS matrices. This makes it numerically more challenging to simulate than the leg direction pumps, even those that inject greater energy. The figure shows that even $\chi=500$ achieves good quantitative agreement for much of the simulation, but does develop an oscillatory error at late times. The largest bond dimensions are all in excellent agreement for almost all times shown.
Appendix B Single time integral approximation
@@
×$\quad\quad\times C(\mathbf{k},\bar{t}+\tau/2,\bar{t}-\tau/2)$
@@
Figure 16: Comparison between the full expression Eq. 4 and the approximation Eq. 6 for the dynamical charge response of a 4 rung ladder using exact diagonalisation, for leg direction pumping with $A_{0}=1$.
Appendix C Effective temperature
@@
Figure 17: The temperature $\beta^{-1}$, extracted using Eq. 17 for the $A_{0}=1$, leg direction pump. Left panel: results using the spin response at two different wave vectors, $(\pi,\pi)$ and $(\pi,0)$. Right panel: the corresponding results from the charge response.
Figure 18: The temperature $\beta^{-1}$, extracted using Eq. 17 for the $A_{0}=1$, rung direction pump. Left panel: results using the spin response at two different wave vectors, $(\pi,\pi)$ and $(\pi,0)$. Right panel: the corresponding results from the charge response.

### v3

@@ IV.3 Charge dynamical structure factor
In contrast to the spin sector, the emergence due to pumping of an intraband charge response at low energies $\omega\sim 0$ is unambiguous because of the large Mott gap. After pumping the intraband charge response is a strong feature, with more spectral weight than the interband charge response in the original equilibrium system. The intraband response arises due to scattering from pump induced doublons. The lowest panels of Fig. 11 and Fig. 12 show the results for the rung pump, which produces fewer doublons than the two leg-direction pumps. Accordingly, the rung pump results show a relatively weaker intraband response (due to fewer photoinduced doublons) and stronger interband response (reflecting a larger density of states for the creation of new doublons by a probe). For the rung-direction pump the intraband response has a pronounced dispersion, becoming gapless on approaching $\mathbf{k}=(0,0)$. This can be interpreted as an effective dispersion for the propagating photoinduced doublons. The leg-direction pumps produce more doublons and the resulting intraband response is flatter: as doublons are hardcore they interact strongly and impede each other, renormalizing the effective charge hopping parameters and reducing the bandwidth.
@@
Figure 13: Charge correlations with $\mathbf{k}=(\pi,\pi)$, at $t=10$ after pumping along the legs $(\parallel)$ or rungs $(\perp)$.
@@
Figure 14: Charge correlations with $\mathbf{k}=(0,\pi)$, at $t=10$ after pumping along the legs $(\parallel)$ or rungs $(\perp)$.
@@
As in the case of the spin response, the detailed effects of the different pumps in the charge sector are best seen by considering cuts at fixed wave vectors. Figs. 13 and 14 show cuts through the charge response at fixed $\mathbf{k}$. Similar to the spin dynamics, the equilibrium charge response at both $(\pi,\pi)$ and at $(\pi,0)$ appears as a resolution limited peak (except for a small feature close to $\omega=6$ for $k_{y}=\pi$). Pumping suppresses this peak and makes it asymmetric, as well as producing a strong low-energy, intraband response. The energetically similar $A_{0}=0.45$ leg-direction, and $A_{0}=1$ rung-direction pumps lead to similar suppression of the interband response at $\omega\sim 10$ for $(\pi,\pi)$ (see Fig. 13). However, there is pump anisotropy visible in the intraband response that develops. Figs. 13 and 14 show that the rung pump produces a weaker response at $\mathbf{k}=(\pi,0)$ than at $\mathbf{k}=(\pi,\pi)$ with an asymmetric energy profile that differs qualitatively between the two wave vectors. In comparison, the $A_{0}=0.45$ leg-direction pump produces a very similar post pump response in terms of intensity and shape at both $k_{y}=\pi$ and $k_{y}=0$.
IV.4 Comparison of numerics with the sum rule
@@
We finish this section by checking that the integrated spin and charge responses obey Eq. 12. Table 2 displays the results of evaluating $W_{s}$ and $W_{c}$ numerically using two different methods, at three different times: t=10t=-10 before the pump, $t=10$ after the pump, and $t=\pi/2$ half way through the pump. The first method involves evaluating Eq. 9 and Eq. 11 using iTEBD data for $\langle S^{2}_{i,l}(t)\rangle$ and $\langle d_{i,l}(t)\rangle$. This method confirms the accuracy of the iTEBD results by satisfying the sum rule $4W_{s}+W_{c}=2G\rho$ to high precision ($10^{-9}$ missing weight).
@@
The second method we use is to integrate the dynamical response $I(\mathbf{k},\omega,t)$, determined using TEBD for the 32 site ladder, over $\omega$ and $\mathbf{k}$. The results of this method for $W_{s}$ and $W_{c}$ are expected to differ from those for iTEBD due to finite size effects and errors introduced by the single integral approximation, Eq. 6. The latter can also cause departures from the sum rule. Concomitantly, in Table 2 the worst agreement with the sum rule occurs at $t=\pi/2$ where the single integral approximation is least accurate. However, the largest error, seen for the leg-direction $A_{0}=0.45$ pump, is still small and $4\%$.
@@ V Conclusions
We have studied the nonequilibrium dynamics of the half-filled Hubbard ladder when it is pumped by an idealised optical pulse, using MPS techniques. We find that the antiferromagnetic correlations that dominate the equilibrium state are significantly reduced in the post pump state: equilibrium AF correlations are strongly suppressed at nearest and next nearest neighbour separations relative to the equilibrium state, before exponential decay sets in with a slightly reduced correlation length. Charge correlations in the initial, equilibrium, state are very weak, but they are strongly enhanced by the pump until they rival the magnetic correlations. We have shown that the anisotropic nature of the ladder means that the energy imparted by the pump is dependent on the pump direction. In general, pumping with a polarisation that alters the hopping parameter along the legs of the ladder leads to greater energy transfer and disruption of spin correlations than pumping polarised in the rung-direction.
A small $t$ expansion and the numerical results show that at short times the response of the system to the pump scales with the number of nearest neighbours and hopping parameters as $\sim z_{\alpha}t_{\alpha}A_{0}^{2}$, where $\alpha=\parallel,\perp$ labels the pump direction. The small $t$ expansion does not rely on the ladder geometry. Hence, our results suggest this scaling will also apply to the short time behaviour of 2D (or 3D) systems. Once $t$ exceeds the characteristic hopping time $t_{h}^{-1}$ the scaling breaks down, and the response of the system is instead more specifically dependent on the pump characteristics (for example, pumps in different directions that have almost equal total photoinjected energies nonetheless produce different numbers of doublon excitations).
@@
We interpret our results in terms of the hard-core nature of the doublon charge excitations of the system. Photodoping excites electrons from the lower to upper Hubbard band, creating doublons. As they are hard-core, each doublon reduces the available phase space for the creation of new doublons, and this leads to a reduced interband response at energies $\omega\sim U$. The photodoping induced doublons lead to a new low energy response due to intraband transitions. This response is sharply defined and dispersing. In spin systems the corresponding intraband behaviour is known as a Villain mode [61]. A similar intraband charge response has been noted in the equilibrium Hubbard chain at temperatures that are finite, but small compared to the Mott gap [62]. A consequence of the strongly correlated nature of excitations in gapped, hard-core, spin systems is that the thermal broadening of the interband response is strongly asymmetric [59, 31, 32]. We see similar asymmetry in the interband charge response after pumping.
@@
We have not found a robust value of $\beta^{-1}$ that is shared for both spin and charge structure factors and at different wave vectors (for details see appendix E). We conclude that, like the chain, the post pump state cannot be simply described using a thermal distribution. Indeed Ref. [38] has used a generalised Gibbs ensemble (GGE) approach to characterise the optically pumped Hubbard chain at late times, with the conservation of doublon number leading to an additional generalised temperature. Checking if the post pump Hubbard ladder can be described using a GGE would be an interesting subject for future work.
@@ Acknowledgements.
The authors thank Andreas Weichselbaum, Mark Dean, Calum MacCormick and Chris Hooley for useful discussions. RMK was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, under Contract No. DE-SC0012704. This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/L010623/1].
Appendix A Short time expansion
The directional dependence of a optical pump on the ladder system can be demonstrated by considering a short time expansion. The ladder Hamiltonian in the absence of pumping can be written as
$H_{0}$ $=H_{\text{leg}}+H_{\text{rung}}+H_{U},$
$H_{\text{leg}}$ $=\sum_{i\ell\sigma}\left[-t_{\parallel}\left(c^{\dagger}_{i\ell\sigma}c_{i+1\ell\sigma}+c^{\dagger}_{i+1\ell\sigma}c_{i\ell\sigma}\right)\right],$
$H_{\text{rung}}$ $=\sum_{i\sigma}\left[-t_{\perp}\left(c^{\dagger}_{i,1,\sigma}c_{i,2,\sigma}+c^{\dagger}_{i,2,\sigma}c_{i,1,\sigma}\right)\right],$
$H_{U}$ $=\sum_{i\ell}n_{i\ell\uparrow}n_{i\ell\downarrow}.$ (14)
The pump induces time-dependent hoppings in the leg and rung-directions,
$H_{\text{leg}}(t)$ $=\sum_{i\ell\sigma}\left[-t_{\parallel}\left(e^{\text{i}A_{\parallel}}c^{\dagger}_{i\ell\sigma}c_{i+1\ell\sigma}+e^{-\text{i}A_{\parallel}}c^{\dagger}_{i+1\ell\sigma}c_{i\ell\sigma}\right)\right],$
$H_{\text{rung}}(t)$ $=\sum_{i\sigma}\left[-t_{\perp}\left(e^{\text{i}A_{\perp}}c^{\dagger}_{i,1,\sigma}c_{i,2,\sigma}+e^{-\text{i}A_{\perp}}c^{\dagger}_{i,2,\sigma}c_{i,1,\sigma}\right)\right],$ (15)
where $A_{\parallel}$ and $A_{\perp}$ are the time-dependent phases given by Peierls substitution. The full time-dependent Hamiltonian can then be written as
$H(t)=H_{\alpha}(t)+H_{0},$ (16)
where for a parallel (leg-direction) pump $\alpha=\parallel$ and
$H_{\parallel}(t)=H_{\text{leg}}(t)-H_{\text{leg}},$ (17)
and for a perpendicular (rung-direction) pump $\alpha=\perp$,
$H_{\perp}(t)=H_{\text{rung}}(t)-H_{\text{rung}}.$ (18)
Consider a pump applied for a very short time $\delta t$, so that the time evolution operator $U(t)$ is given approximately by
$U(t)\approx\exp[-\text{i}H(\delta t)\delta t].$ (19)
We assume that the pump is switched on suddenly at $t=0$ (as in our numerics) but suppressed by a Gaussian factor, so that $A_{\alpha}(\delta t)\approx A_{0}\exp[-t_{p}^{2}/(2\sigma_{p}^{2})]\Omega\delta t\ll 1$.
The energy that is imparted by the pump in this short time $\delta t$, relative to the energy of the initial state, is
$\delta E=\langle e^{\text{i}H(\delta t)\delta t}H_{0}e^{-\text{i}H(\delta t)\delta t}\rangle-\langle H_{0}\rangle$ (20)
where the expectation is evaluated with respect to the ground state $|\psi_{0}\rangle$ of the initial Hamiltonian. Expanding the exponentials gives
$\delta E=$ $\left\langle\text{i}\left[H(\delta t),H_{0}\right]\delta t-\frac{1}{2}\left\{H^{2}(\delta t),H_{0}\right\}\delta t^{2}\right$.
[…]
The commutator in the first term vanishes because the initial state is an eigenstate of $H_{0}$, that is, $H_{0}|\psi_{0}\rangle=E_{0}|\psi_{0}\rangle$. Writing $H=H_{0}+H_{\alpha}$, expanding and then collecting terms gives
$\delta E$ $=\langle H_{\alpha}(\delta t)(H_{0}-E_{0})H_{\alpha}(\delta t)\rangle\delta t^{2}+O(\delta t^{3})$ (22)
Further explicit evaluation would require knowledge of the precise form of $|\psi_{0}\rangle$. However, the dependence of $\delta E$ can be extracted heuristically as follows. Processes that contribute to Equation 22 involve a hop with coefficient $t_{\alpha}(e^{\pm\text{i}A_{\alpha}}-1)\approx\pm\text{i}t_{\alpha}A_{\alpha}$, a double occupancy penalty, $H_{0}-E_{0}\sim U$ and then a hop back $t_{\alpha}(e^{\mp\text{i}A_{\alpha}}-1)\approx\mp\text{i}t_{\alpha}A_{\alpha}$. For leg-direction quenches this gives a contribution $\sim 2t_{\parallel}A_{\parallel}^{2}U$ per site where the factor of two arises due to the two nearest neighbours in the leg-direction. For rung-direction quenches the contribution is $\sim t_{\perp}A_{\perp}^{2}U$ per site, because each site only has one nearest neighbour in the rung-direction. Hence, the anisotropic nature of the Hamiltonian results in an energy dependence on pump direction even at the shortest times, and when $t_{\perp}=t_{\parallel}$ energy transfer to the ladder is enhanced for pumps in the leg-direction relative to the rung-direction.
Appendix B Pump frequency
The results in the main body of this work were obtained using a pump frequency $\Omega=6$. This value was chosen to enable comparison with a previous nonequilibrium investigation of the Hubbard chain. In this appendix we consider other values of $\Omega$ and their effect on the injected energy density.
Figure 15: The change in energy density after the pump, as a function of the pump frequency $\Omega$ for $A_{0}=1$. Results are shown for a finite ladder with $L=8$ rungs and periodic boundary conditions (pbc). The infinite system results for $\Omega=6$ are also plotted for comparison.
The response of the ladder to pumping depends on the pump frequency $\Omega$. For very low frequencies, $\Omega\ll U$, the response is almost adiabatic: the state of the system will closely track the instantaneous ground state of the time-dependent Hamiltonian and at the end of the pump the system will be in a state that closely approximates the initial state, so $\Delta E\approx 0$. At high frequencies, when $\Omega$ exceeds the inverse electron hopping timescale, the electrons cannot react fast enough to respond to the pump. This limits energy transfer and $\Delta E$ again approaches zero. Between these extremes, a strong response is expected for frequencies $\Omega\sim U$, close to the Mott gap.
Fig. 15 shows the injected energy density as a function of pump (angular) frequency $\Omega$ for an infinite ladder and for an 8 rung ladder with periodic boundary conditions. The form of the pump is given by Eq. 2 with amplitude parameter $A_{0}=1$ and width $\sigma_{p}=1$.
For both pumping directions $\Delta E$ increases with frequency, reaching a maximum at $\Omega\approx U$ ($\Omega=10$ for the leg-direction and $\Omega=8$ for the rung-direction) before declining. This is consistent with the expectation that pumping with a frequency near the Mott gap should yield the strongest response.
Appendix C Effect of bond dimension
@@
Our simulations are carried out using infinite and finite MPS with bond dimensions of up to $\chi=3000$, and we perform time evolution using iTEBD and TEBD with a second order Trotter decomposition and time step, $\Delta t=0.01$. In general we find truncation errors of order $10^{-6}$ at the latest times we report, but because correlations become very short ranged after the pump this does not significantly affect our results (especially once the experimental resolution $\sigma_{X}$ is taken into account). As an example, in Fig. 16 we plot $\langle d_{i,l}\rangle$ for the rung-direction pump with $A_{0}=1.0$ using iTEBD with different bond dimensions, $\chi$.
@@
Figure 16: Doublon and holon density as the Hubbard ladder is pumped in the rung-direction, i.e. $(\perp,A_{0}=1.0)$, at different bond dimensions $\chi$.
@@
The rung pump case breaks parity symmetry between the legs leading to a less compact structure for the MPS matrices. This makes it numerically more challenging to simulate than the leg-direction pumps, even those that inject greater energy. The figure shows that even $\chi=500$ achieves good quantitative agreement for much of the simulation, but does develop an oscillatory error at late times. The largest bond dimensions are all in excellent agreement for almost all times shown.
Appendix D Single time integral approximation
@@
$\quad\quad\times C(\mathbf{k},\bar{t}+\tau/2,\bar{t}-\tau/2)$
@@
Figure 17: Comparison between the full expression Eq. 4 and the approximation Eq. 6 for the dynamical charge response of a 4 rung ladder using exact diagonalisation, for leg-direction pumping with $A_{0}=1$.
Appendix E Effective temperature
@@
Figure 18: The temperature $\beta^{-1}$, extracted using Eq. 26 for the $A_{0}=1$, leg-direction pump. Left panel: results using the spin response at two different wave vectors, $(\pi,\pi)$ and $(\pi,0)$. Right panel: the corresponding results from the charge response.
Figure 19: The temperature $\beta^{-1}$, extracted using Eq. 26 for the $A_{0}=1$, rung-direction pump. Left panel: results using the spin response at two different wave vectors, $(\pi,\pi)$ and $(\pi,0)$. Right panel: the corresponding results from the charge response.